Chance Company had two operating divisions, one manufacturing farm equipment and the other office supplies. Both divisions are considered separate components as defined by generally accepted accounting principles. The farm equipment component had been unprofitable, and on September 1, 2018, the company adopted a plan to sell the assets of the division. The actual sale was completed on December 15, 2018, at a price of $680,000. The book value of the division’s assets was $1,170,000, resulting in a before-tax loss of $490,000 on the sale.

The division incurred a before-tax operating loss from operations of $180,000 from the beginning of the year through December 15. The income tax rate is 40%. Chance’s after-tax income from its continuing operations is $450,000.

Required:
Prepare an income statement for 2018 beginning with income from continuing operations. Include appropriate EPS disclosures assuming that 100,000 shares of common stock were outstanding throughout the year. (Amounts to be deducted should be indicated with a minus sign. Round EPS answers to 2 decimal places.)

CHANCE COMPANY
Partial Income Statement
For the Year Ended December 31, 2018
Income from continuing operations
Discontinued operations:
0
$0
Earnings per share:
$0.00
Answer #1
Chance Company
Partial Income Statement
For the Year ended December 31st
Income from continuing operations=(A) $   4,50,000.00
Discontinued operations gain/loss
Loss from operations of discontinued component($490000+$180000) $ -6,70,000.00
Income tax benefit=($670000*40%) $   2,68,000.00
Loss on discontinued operations=(B) $ -4,02,000.00
Net Income=(A)-(B) $       48,000.00
Earnings per share
Income from continued operations=(A) $   4,50,000.00
Number of shares outstanding=(B) 100000
EPS=(A)/(B)=(C ) $                 4.50
Loss from discontinued operations=(A) $   4,02,000.00
Number of shares outstanding=(B) 100000
EPS=(A)/(B)=(D) $                 4.02
Net Income=(C )-(D) $                 0.48